Consider a company with total dividends of $0.5 per share for the mmost recent year
(year=0). The company's dividends are expected to grow at a rate of 10% for the next 3
years and at a constant rate of 5% thereafter. Given the risk of the business, the required
rate of return is 15%. What is the value of the firm's equity?

Please show your work. Thank you.
Answer #1

Last Dividend, D0 = $0.50

Growth rate for next 3 years is 10% and a constant growth rate (g) of 5% thereafter

D1 = $0.50 * 1.10 = $0.55
D2 = $0.55 * 1.10 = $0.605
D3 = $0.605 * 1.10 = $0.6655
D4 = $0.6655 * 1.05 = $0.698775

Required Return, rs = 15%

P3 = D4 / (rs - g)
P3 = $0.698775 / (0.15 - 0.05)
P3 = $0.698775 / 0.10
P3 = $6.98775

P0 = $0.55/1.15 + $0.605/1.15^2 + $0.6655/1.15^3 + $6.98775/1.15^3
P0 = $5.97

So, value of the firm’s equity is $5.97